OpenStudy (anonymous):

Solve each system of equations by using elimination. a. 3.19, 0 c. 1.57, –0.81 b. 3.19, –0.81 d. 1.38, 3.19

OpenStudy (anonymous):

where is the system of equations ?

OpenStudy (anonymous):

8p+8q=19 5p-5q=20 @kelliegirl33

OpenStudy (anonymous):

8p + 8q = 19 -->(-5) 5p - 5q = 20 --> (8) ------------------- -40p - 40q = - 95 (result of multiplying by -5) 40p - 40q = 160 (result of multiplying by 8) -------------------add -80q = 65 q = -65/80 q = - 0.8125 or - 0.81 now sub -0.81 in for q in any of the beginning equations 5p - 5q = 20 5p - 5(-0.81) = 20 5p + 4.05 = 20 5p = 20 - 4.05 5p = 15.95 p = 15.95/5 p = 3.19 check... 8p + 8q = 19 8(3.19) + 8(-0.81) = 19 25.52 - 6.48 = 19 19.04 rounds to 19 = 19 (correct) p = 3.19 q = 0.81